# TinyChemVL: Advancing Chemical Vision-Language Models via Efficient Visual Token Reduction and Complex Reaction Tasks

### v1

@@ Abstract
While Vision Language Models (VLMs) have demonstrated remarkable capabilities in general visual understanding, their application in the chemical domain has been limited, with previous works predominantly focusing on text and thus overlooking critical visual information, such as molecular structures. Current approaches that directly adopt standard VLMs for chemical tasks suffer from two primary issues: (i) computational inefficiency of processing entire chemical images with non-informative backgrounds. (ii) a narrow scope on molecular-level tasks that restricts progress in chemical reasoning. In this work, we propose TinyChemVL, an efficient and powerful chemical VLM that leverages visual token reduction and reaction-level tasks to improve model efficiency and reasoning capacity. Also, we propose ChemRxn-V, a reaction-level benchmark for assessing vision-based reaction recognition and prediction tasks. Directly predicting reaction products from molecular images poses a non-trivial challenge, as it requires models to integrate both recognition and reasoning capacities. Our results demonstrate that with only 4B parameters, TinyChemVL achieves superior performance on both molecular and reaction tasks while demonstrating faster inference and training speeds compared to existing models. Notably, TinyChemVL outperforms ChemVLM while utilizing only 1/16th of the visual tokens. This work builds efficient yet powerful VLMs for chemical domains by co-designing model architecture and task complexity.
@@ Introduction
The remarkable success of Large Language Models (LLMs) on diverse tasks, from question answering to code generation (zhang2024internlm; grattafiori2024llama), has spurred significant interest in leveraging their reasoning capabilities to accelerate scientific discovery. This trend is particularly prominent in domains like mathematics and chemistry, where core entities such as equations and molecules can be parsed in structured textual formats (lu2022dynamic; lightman2023let; zhang2024chemllm; zhang2025scientific). Especially in the chemical field, the molecules are parsed into SMILES (lightman2023let) and SELFIES (krenn2022selfies) formats for the following analysis. However, representing the chemical molecules in text format inevitably omits some spatial information. More recently, Vision Language Models (VLMs) have demonstrated remarkable capabilities in addressing a wide range of visual tasks, such as image captioning and question answering (liu2023llava; zhao2025chartcoder), which can handle more complex tasks. However, in the chemical field, how to efficiently utilize VLMs on chemical tasks still lacks exploration. Recent works have demonstrated initial attempts in adapting VLMs for chemical applications such as molecule caption and property prediction (li2025chemvlm). However, the models are generally fine-tuned on general domain VLMs, which limits their efficiency in handling chemical problems (bi2025prism).
@@
Recently, leveraging the projection layer to align LLMs and vision encoders, VLMs have achieved superior performance on many visual tasks (zhang2024mm; bi2024visual). Current models, including GPT-4o (openai2024gpt4o), InternVL (chen2024expanding) and Qwen-VL (bai2023qwenvlversatilevisionlanguagemodel) have shown strong performance on vision understanding and reasoning tasks. Recently, VLMs have been increasingly applied to vision-based tasks in the field of chemistry. For instance, ChemVLM (li2025chemvlm) introduces vision-based chemical datasets and utilizes InternVL architecture (chen2024expanding) as the backbone for fine-tuning. ChemDFM-X (zhao2024chemdfmx) utilizes distinct encoders to align representations from visual and graph modalities with the texts. In the generative domain. ChemMLLM (zhang2024chemllm) employs the VQ-GAN (esser2021taming) as the encoder and decoder for molecular image understanding and generation. However, these models are typically directly fine-tuned on existing architectures and suffer from significant visual token redundancy, resulting in computational inefficiency. To solve these problems, we propose TinyChemVL, which adopts an adaptive token merging and pruning strategy and a large-scale and diverse training corpus to improve both model efficiency and performance.
@@ Methods
Figure 1: The overview structure of TinyChemVL. For the input image, a dynamic resolution strategy is employed to segment and downsample the input image to standard resolutions. Subsequently, the vision encoder progressively reduces the number of visual tokens through the adaptive token merging and pruning method. The pixel shuffle operation is the default downsampling strategy utilized in InternVL2.5.

### v2

@@ Abstract
While Vision Language Models (VLMs) have demonstrated remarkable capabilities in general visual understanding, their application in the chemical domain has been limited, with previous works predominantly focusing on text and thus overlooking critical visual information, such as molecular structures. Current approaches that directly adopt standard VLMs for chemical tasks suffer from two primary issues: (i) computational inefficiency of processing entire chemical images with non-informative backgrounds. (ii) a narrow scope on molecular-level tasks that restricts progress in chemical reasoning. In this work, we propose TinyChemVL, an efficient and powerful chemical VLM that leverages visual token reduction and reaction-level tasks to improve model efficiency and reasoning capacity. Also, we propose ChemRxn-V, a reaction-level benchmark for assessing vision-based reaction recognition and prediction tasks. Directly predicting reaction products from molecular images poses a non-trivial challenge, as it requires models to integrate both recognition and reasoning capacities. Our results demonstrate that, with only 4B parameters, TinyChemVL achieves superior performance on both molecular and reaction tasks, while also demonstrating faster inference and training speeds compared to existing models. Notably, TinyChemVL outperforms ChemVLM while utilizing only 1/16th of the visual tokens. This work builds efficient yet powerful VLMs for chemical domains by co-designing model architecture and task complexity.
@@ Introduction
The remarkable success of Large Language Models (LLMs) on diverse tasks, from question answering to code generation (zhang2024internlm; grattafiori2024llama), has spurred significant interest in leveraging their reasoning capabilities to accelerate scientific discovery. This trend is particularly prominent in domains like mathematics and chemistry, where core entities such as equations and molecules can be parsed in structured textual formats (lu2022dynamic; lightman2023let; zhang2024chemllm; zhang2025scientific). Especially in the chemical field, the molecules are parsed into SMILES (lightman2023let) and SELFIES (krenn2022selfies) formats for the following analysis. However, representing the chemical molecules in text format inevitably omits some spatial information. More recently, Vision Language Models (VLMs) have demonstrated remarkable capabilities in addressing a wide range of visual tasks, such as image captioning and question answering (liu2023llava; zhao2025chartcoder), which can handle more complex tasks. However, in the chemical field, there is still a lack of exploration on how to efficiently utilise VLMs for chemical tasks. Recent works have demonstrated initial attempts in adapting VLMs for chemical applications such as molecule caption and property prediction (li2025chemvlm). However, the models are generally fine-tuned on general domain VLMs, which limits their efficiency in handling chemical problems.
@@
Recently, leveraging the projection layer to align LLMs and vision encoders, VLMs have achieved superior performance on many visual tasks (zhang2024mm; bi2024visual). Current models, like GPT-4o (openai2024gpt4o) and InternVL (chen2024expanding) have shown strong performance on vision understanding and reasoning tasks. Recently, VLMs have been increasingly applied to vision-based tasks in the field of chemistry. For instance, ChemVLM (li2025chemvlm) introduces vision-based chemical datasets and utilizes InternVL architecture (chen2024expanding) as the backbone for fine-tuning. ChemDFM-X (zhao2024chemdfmx) utilizes distinct encoders to align representations from visual and graph modalities with the texts. In the generative domain. ChemMLLM (zhang2024chemllm) employs the VQ-GAN (esser2021taming) as the encoder and decoder for molecular image understanding and generation. However, these models are typically directly fine-tuned on existing architectures and suffer from significant visual token redundancy, resulting in computational inefficiency. To address these issues, we propose TinyChemVL, which employs an adaptive token merging and pruning strategy, along with a large-scale and diverse training corpus, to enhance both model efficiency and performance.
@@ Methods
Figure 1: The overview structure of TinyChemVL. For the input image, a dynamic resolution strategy is employed to segment and downsample it to standard resolutions. The vision encoder progressively reduces the number of visual tokens through the adaptive token merging and pruning method. The pixel shuffle operation is the default strategy utilized in InternVL2.5.
@@ Conclusion
Acknowledgements
This work was supported by the robotic AI-Scientist platform of Chinese Academy of Sciences.